From GlueXWiki
Jump to: navigation, search

May 10, 2012 FDC meeting


  1. Production Construction Tracking (Dave)
    • Production status
    • Third package refurbishment status
    • Third package tests (Lubomir)
  2. Engineering update
  3. Electronics update (Chris)
  4. Tests with first package in 126 FDC E-log (Beni)
  5. Other


Participants: Dave, Chris, Nick, Simon, Beni, Mark, and Lubomir.


- Dave for package #4 production: working now on four cathodes, there's one more and we will be done with all the cathodes for this package, then continuing with the spare cathodes. Chris will put the elements on wire plane #26; #24 and #25 were deadened. Wire frame #27 is on the stringing table. Received wire frame #18 back from the machine shop. They fixed some of the frame holes that were not correct, these are used to attach the frame to the strong-back for stringing.

- Today we continued with the installation of the last two cells of the third package. Before the installation, in one of the new wire planes we found one broken wire (between the pad and the epoxy) that was fixed on Tuesday and checked on Wednesday. After the meeting all the cells of package #3 were installed and started flushing with gas.

- All the four cells of the third package, as tested with a source, are OK. The oxygen for the four cells went down to 100ppm in about a week. We had some problems applying HV on the bottom cell (that was tested already before when it was installed alone) but we were able to condition it and was stable for >48 hours.


- This morning we had issues with the new installation pins that we used to replace the old that were bent. The tolerance is only 0.5mills! We suspect that one of the hub (aluminum) was damaged by the steel pin. Bill suggested fixing it after we are done with the third package. In the mean time Anatoly suggested putting vacuum grease (same that we use on the O-rings) on the pin when inserting it in the hub and the package. This worked very well and we were able to use all the pins. The problem with the pins is that they can easily be bent when tightening the handle. Casey also suggested to put two handles, one at the bottom of the pin and the other to leave it loose and tighten it at the right place at the end before tightening the package. This worked also very well.


- For the grounding Chris got an offer for 300 clips, 100 of each type. Beni explained that we need only one type: 11mm width. Chris will ask the company again. The thickness of the material to be defined experimentally when we get some samples. We will need 20 clips per cell, or ~120 for one package.

Tests with first package at 126

- Beni: HV now is 1600V to avoid afterpulsing, the efficiency is ~90% except two cells with areas below 80%. The position resolution is still of the order of 300 microns. Beni explained the procedure he uses to estimate the resolution: the incident angle from the external tracking and then comparing the positions from parallel wires the pairs of cells that have wires with the same direction. We discussed different sources that could make the resolution worse: how parallel the wires on the different cells are, the time to distance conversion in presence of oxygen and low efficiency for big drift times. Also Beni wants to look in the external chambers showing some uneven distributions of the cosmics.